anton is four years older than liza. the product of his age 3 years from now and liza's age two years ago is 90 years old. how old they are?
Answer by jorel555(1290)

Let l be Liza. Then Anton is l+4. So:
(l+4+3)(l-2)=90
l^2+5l-104=0
(l+13)(l-8)=0
l=-13, 8
Liza has to be 8 years old, and Anton is 12.😊😊😊😊